OpenStudy (anonymous):

The area of a triangle is 12 square inches. The base is 2 inches longer than the height. Find the length of the base and the height using quadratic formula.

OpenStudy (anonymous):

A= 1/2 *h* b right?

OpenStudy (anonymous):

hm so.... 12 = h^2* 2h * 1/2 ? right?

OpenStudy (anonymous):

Hello, sorry nobody answered your question. I'm just going to do the work and see if I end up with the same quadratic you did. The height is two inches larger than the base. So height = base + 2 inches. 12 = (1/2)(height)(base) 12 = (1/2)(base+2)(base) 12 = (1/2)(base^2 + 2base) 0 = (1/2)(base^2) + (1)base - 12 0 = (1/2)(b^2) + b - 12

OpenStudy (anonymous):

From here, you'd use the quadratic equation to find both roots. If I can help you through the steps I went through, let me know.

OpenStudy (dumbcow):

this can also be done by factoring 1/2(b^2 +2b) = 12 b^2 + 2b = 24 b^2 +2b-24 = 0 (b+6)(b-4) = 0 b = -6, b = 4 ignore the neg answer b=4 and b+2 = 6